OpenStudy (anonymous):

Evaluate log(small)9 (big)27.

OpenStudy (agreene):

\[\log_9 (27)=\frac{\ln(27)}{\ln(9)}=\frac{3}{2}==1.5\]

sam (.sam.):

\[\log_{9}27=\frac{\log_{3}27}{\log_{3}9}=\frac{\log_{3}3^{3}}{\log_{3}3^{2}}=\frac{3}{2}\]